OpenStudy (anonymous):

Let f(x) = 1/6 x + 9 , \quad 2 \le x \le 9 The domain of f^(-1) is the interval [A,B]

OpenStudy (anonymous):

\[f(x) = 1/6 x + 9 , \quad 2 \le x \le 9 \] The domain of f^(-1) is the interval [A,B] Where A= B=